### v2

# A novel sensitivity analysis method for agent-based models stratifies in-silico tumor spheroid simulations

Abstract
Agent-based models (ABMs) are widely used in biology to understand how individual actions scale into emergent population behavior. Modelers employ sensitivity analysis (SA) algorithms to quantify input parameters’ impact on model outputs, however, it is hard to perform SA for ABMs due to their computational and complex nature. In this work, we develop the Simulate, Summarize, Reduce, Cluster, and Analyze (SSRCA) methodology, a machine-learning based pipeline designed to facilitate SA for ABMs. In particular, SSRCA can achieve the following tasks for ABMS: 1) identify sensitive model parameters, 2) reveal common output model patterns, and 3) determine which input parameter values generate these patterns. We use an example ABM of tumor spheroid growth to showcase how SSRCA provides similar SA results to the popular Sobol’ Method while also identifying four common patterns from the ABM and the parameter regions that generate these outputs. This analysis could streamline data-driven tasks, such as parameter estimation, for ABMs by reducing parameter space. While we highlight these results with an ABM on tumor spheroid formation, the SSRCA methodology is broadly applicable to biological ABMs.
1 Introduction
Multicellular tumor spheroids are a widely-used experimental procedure to study avascular tumor growth and anti-cancer drug therapies [1, 2, 3]. As opposed to 2-dimensional assay experiments, 3-dimensional tumor spheroid experiments closely mimic the in vivo tumor microenvironment with realistic cell-cell interactions, chemical gradients, the formation of hypoxic regions, etc. These conditions lead to high levels of intratumoral heterogeneity. For example, cells in oxygen rich environments will rapidly divide and increase in number, whereas those in hypoxic regions may become quiescent or die. The fluorescent ubiquitination-based cell cycle indicator (FUCCI) imaging technique is a useful modality to measure and quantify tumor spheroid heterogeneity by tracking individual cells’ progression through the cell cycle in real-time [3, 4, 5]. Using FUCCI, the nuclei of cells in G1 phase fluoresce red, while the nuclei of cells in G2/S/M phase fluoresce green. The nuclei of cells in early S (eS) phase fluoresce yellow due to the presence of both red and green fluorescence [4].
Mathematical models provide an efficient and cost-effective approach to generate hypotheses for tumor spheroid formation and optimally design future experiments [6, 7]. Common approaches for simulating in-silico tumor spheroids include continuum models [8, 9, 10, 11, 12, 13] and cellular automata [14, 15, 16]. Agent-based models (ABMs) are another popular modeling choice due to their ability to capture the discrete and stochastic nature of tumor spheroids [17, 18, 19, 14, 20]. In an ABM, modelers simulate pre-defined individual-level rules to determine how they translate into collective behavior. For example, Klowss et al. developed a hybrid ABM that consists of both discrete tumor cells and a continuous variable representing a nutrient (such as oxygen) that controls cells’ progression through the cell cycle [21]. This model includes rules dictating cell migration, death, and transition between cell cycle states. The resulting in silico tumor spheroid captures many realistic facets of in vitro tumor spheroid growth: an outer proliferating shell, an inner growth-arrested shell, and a necrotic core [22]. Model simulations serve to precisely predict how experimental conditions and/or design choices (including initial spheroid size, kinetic parameter rates, time duration, etc.) impact these patterns. In another study, Rocha et al. demonstrated that their epithelial metastasis ABM could be used as a cancer patient digital twin to develop patient-specific models for precision medicine treatment plans [23].
The analysis of ABMs presents many challenges; notably, ABMs often require many parameter values, are not amenable to extensive simulation due to long simulation times, and their outputs are complex and noisy. These difficulties hinder the application of standard modeling techniques for ABMs, including parameter estimation, uncertainty quantification, and sensitivity analysis. As one example, consider the 3-dimensional Klowss Model from [21]: There are 25 model parameters: 13 can be estimated from biological experiments, 2 can be determined from numerical experiments, leaving 10 unknown parameters. Impressively, the authors selected specific values for all 10 parameters and demonstrated excellent agreement between their in-silico simulations and in-vitro spheroid experiments using human primary melanoma cells. However, there is no guarantee that the same values would generalize to data from other cell lines. Estimating all 10 parameter values from new data would require exhaustively searching the 10-dimensional parameter space, which is infeasible due to long simulation times and the curse of dimensionality [24]. In addition to parameter estimation, these obstacles inhibit our ability to perform feature mapping, which links the model’s input parameter values to its output patterns. In the context of tumor spheroids, these limitations prevent the use of complex models for precision medicine diagnostics [25].
Parameter sensitivity analysis (SA) is an area of research to identify a model’s most influential parameters. Many aspects of the modeling process are encompassed in SA, including quantifying the uncertainty of model outputs, determining influential input parameters, and feature mapping [26, 27]. Modelers can fix the insensitive parameters to reduce the dimensionality of the input parameter space while still capturing the main model behavior. There are many different approaches to performing SA [28, 29]. For example, the Morris one-at-a-time (MOAT) method is a computationally-efficient algorithm that varies one parameter at a time. This simple approach ignores interactions between parameters and can thus generate misleading results for complex models. The Sobol’ Method is a global SA (GSA) approach that ranks parameters based on their contribution to the model’s variance using computed sensitivity index values. These Sobol’ index values incorporate parameter interactions, but this added information comes at the cost of higher computational expenses. Regression-based methods, such as partial rank correlation coefficient, are another popular SA method in which one fits a regression model to predict the model’s output values from its input parameters. These approaches can be used to perform feature mapping, however, it can be challenging to determine a suitable regression model for nonlinear models [30].
Despite the wide and successful use of SA methods, their application to ABMs has been limited. For example, feature mapping is challenging to perform for an ABM because nearby parameter values can yield divergent patterns. Broeke et al. assessed the performance of MOAT and the Sobol’ Method for ABMs, and found that the simple MOAT approach can be used for preliminary feature mapping analysis, but that the Sobol’ Method cannot because its analyses are computed over the entire parameter space instead of considering distinct parameter regions [26]. Other challenges for performing SA for ABMs arise due to their time-varying, nonlinear, computational, and multi-scale natures [31, 27, 30]. Instead of applying standard GSA methods to ABMs, there is an urgent need for the development of novel ABM-specific SA approaches to fully harness the potential of ABMs for modeling complex phenomena.
Towards this end, we introduce the Simulate, Summarize, Reduce, Cluster, and Analyze (SSRCA, pronounced “circa”) methodology to perform GSA for ABMs. An advantage of this approach over other GSA algorithms, such as the Sobol’ indices method, is that it can both identify sensitive parameters and perform feature mapping. In particular, SSRCA can achieve the following tasks: 1) identifies sensitive model parameters, 2) reveals common output model patterns, and 3) determines which parameter values generate these patterns. We focus our work on a 2-dimensional verision of the Klowss Model from [21] using two simulated datasets. For a small dataset (generated by varying 2 model parameters), SSRCA directly links parameter values to one of four typical model patterns. For a larger dataset (generated by varying 10 model parameters), SSRCA classifies four sensitive model parameters, and estimates the marginal parameter distributions underlying each such pattern. These results suggest that SSRCA can describe the parameter values (and, in turn, biological mechanisms) underlying each output model pattern for complex ABMs. This study is focused on a model of tumor spheroid growth, however, the SSRCA methodology is broadly applicable to biological ABMs over all scales, ranging from intracellular protein dynamics to disease spread in human populations [32, 33].
2 Methods
In this section, we review the Klowss model from [21] 2.1, the SSRCA methdology 2.2, and the Sobol’ Method 2.3.
2.1 The Klowss Model of tumor spheroid growth
We implement our proposed methodology on the Klowss Model, an off-lattice hybrid ABM of tumor spheroid growth [21]. We provide details on the model and its implementation in Appendix A. Briefly, the model contains $N(t)$ agents that are colored according to their cell cycle status over time: agents in the G1 phase are colored red, agents in the early S stage are yellow, agents in the S/G2/M phase are green, and dead cells are blue (Figure 1(a)). All agents perform rules on migration, death, and cell cycle progression; when an agent transitions from green to red, it creates an additional red daughter cell to imitate cell division. (Figure 1(b-d)). The kinetic rates for each rule depend on a continuous variable, $c(\bm{x},t)$, such as oxygen (Figure 1(e)). The spatiotemporal concentration of $c(\bm{x},t)$ varies due to diffusion and agent consumption.
Figure 1: Tumor spheroid model overview. (a) Agent cells are colored according to their cell cycle status. (b) All agents migrate with rate $m(c)$, and move the length of a cell diameter in a randomly chosen direction. (c) All agents die with rate $d(c)$ and change their cell cycle status to dead. (d) The cell cycle is modeled as follows: red agents transition to yellow with rate $R_{r}(c)$, yellow agents transition to green with rate $R_{y}(c)$, green agents undergo mitosis with rate $R_{g}(x)$ in which they become red and create a daughter red cell; the daughter cell is placed one cell diameter away in a randomly chosen direction. (e) All model rates depend on the nutrient available at the agent’s physical location.
The Klowss Model from [21] is 3-dimensional in space, however, we find that simulating the model in 2-dimensions can provide good agreement with the middle cross-section of a 3-dimensional simulation. As a computationally-efficient alternative to this 3-dimensional model, we compute the model in 2-dimensions ($\bm{x}_{n}(t)=(x_{n}(t),y_{n}(t))^{T}$) to ease the model’s computational requirements. This is a common simplification used when performing large-scale ABM simulations [23, 34, 35, 36].
2.2 The SSRCA methodology
We introduce the SSRCA methodology, which is designed to perform GSA for ABMs (Figure 2). The SSRCA methodology pipeline includes 5 steps:
Figure 2: The SSRCA methodology pipeline. The SSRCA sensitivity analysis pipeline consists of 5 steps. 1. Simulate model data: Simulate a large number of model simulations from sampled parameter values to generate a dataset. 2. Summarize simulations: Compute uniform length descriptor vectors (DVs) to summarize each model simulation. 3. Reduce dimension: Perform dimensionality reduction to reduce each DV to a dimensionality-reduced DV (DRDV). 4. Cluster data: Cluster all DRDVs into $k$ distinct groups. 5. statistical Analysis: Analyze the clusterings and their constituent parameter samples by a) assessing cluster robustness, b) visually inspecting each cluster’s data, and/or performing statistical tests to infer cluster differences.
  1. 1.
    Simulate model data (Section 2.2.1),
  2. 2.
    Summarize simulations (Section 2.2.2),
  3. 3.
    Reducing dimension (Section 2.2.3),
  4. 4.
    Cluster data (Section 2.2.4), and
  5. 5.
    statistical Analysis (Section 2.2.5).
We now detail all five steps of this methodology. All code and simulated data for this study is implemented in Python (version 3.10.11) and publicly available at https://github.com/e-rohr/FUCCI_ABM2D.
2.2.1 Simulate model data
We simulate the model over a sample of parameter values from a specified distribution. In this study, we apply the SSRCA methodology to two simulated datasets to illustrate its performance on multiple data scales. A brief description of each parameter, including which parameters are varied and the minimum and maximum values of all varied parameters, is provided in Table 1. For simplicity, all parameters were varied between half of their baseline value and twice their baseline value.
Parameter name Symbol Base value Varying Minimum Maximum
maximum death rate $d_{max}$ 2 $\mathrm{h}^{-1}$ yes 1 $\mathrm{h}^{-1}$ 4 $\mathrm{h}^{-1}$
minimum death rate $d_{min}$ 0.0005 $\mathrm{h}^{-1}$ yes 0.00025 $\mathrm{h}^{-1}$ 0.001 $\mathrm{h}^{-1}$
maximum migration rate $m_{max}$ 0.12 $\mathrm{h}^{-1}$ yes 0.06 $\mathrm{h}^{-1}$ 0.24 $\mathrm{h}^{-1}$
minimum migration rate $m_{min}$ 0.06 $\mathrm{h}^{-1}$ yes 0.03 $\mathrm{h}^{-1}$ 0.12 $\mathrm{h}^{-1}$
hill function index for arrest $\eta_{1}$ 5 yes 2.5 10
hill function index for migration $\eta_{2}$ 5 yes 2.5 10
hill function index for death $\eta_{3}$ 15 yes 7.5 30
critical arrest concentration $c_{a}$ 0.4 yes 0.2 0.8
critical migration concentration $c_{m}$ 0.5 yes 0.25 1
critical death concentration $c_{d}$ 0.1 yes 0.05 0.2
initial number of agents $N(0)$ 1100 no N/A N/A
domain length $L$ 1000 $\mu\mathrm{m}$ no N/A N/A
initial spheroid radius $r_{o}(0)$ 245 $\mu\mathrm{m}$ no N/A N/A
dispersal/migration distance $\mu$ 12 $\mu\mathrm{m}$ no N/A N/A
simulation time $T$ 240 $\mathrm{h}$ no N/A N/A
maximum G1-eS transition rate $R_{r}$ 0.047 $\mathrm{h}^{-1}$ no N/A N/A
constant eS-S/G2/M transition rate $R_{y}$ 0.50 $\mathrm{h}^{-1}$ no N/A N/A
[…]
number of nodes $I^{2}$ $200^{2}$ no N/A N/A
steady-state solution interval $t^{*}$ 1 $\mathrm{h}$ no N/A N/A
consumption-diffusion ratio $\alpha$ 0.01 $\mu\mathrm{m}$ cell-1 no N/A N/A
Table 1: ABM parameters. For all the perturbed parameters, the minimum parameter value is always half the baseline value, while the maximum parameter value is always double the baseline value.
Dataset 1: In this dataset, we vary the hill function index for arrest, $\eta_{1}$, and the critical arrest concentration, $c_{a}$, and set all other parameters to their base values from Table 1. We vary both $\eta_{1}$ and $c_{a}$ over 11 logarithmically spaced values that range over their minimum and maximum values, totaling 121 parameter combinations. We perform 10 model simulations for each combination, resulting in $11\times 11\times 10=1,210$ total model simulations.
Dataset 2: In this dataset, we vary two parameters at a time while fixing all others at their base values. We consider all (${10\choose 2}=45$ pairs of the varying parameters from Table 1. For each pair, we vary both parameters over 11 logarithmically spaced values that range over their minimum and maximum values, totaling 121 parameter combinations for each pair of parameters. We perform 10 model simulations for each parameter combination, leading to $45\times 11\times 11\times 10=54,450$ total simulations.
2.2.2 Summarize simulations
Recording the locations and cell cycle states of all $N(t)$ agents over time results in prohibitively large datasets because each simulation contains thousands of agents. Furthermore, standard machine learning algorithms require fixed length vectors, but each simulation has a different number of cell counts. For these reasons, we compute fixed length descriptor vectors (DVs) to concisely summarize each model simulation. We consider two separate DVs in this study: cell subpopulation counts over time, and the final living cell density over space. Other DVs are possible; the appropriate DV choice depends on the model approach and application [34, 36].
[…]
Final Cell Density: We create a two-dimensional histogram of living agent (those colored red, yellow, and green) locations at the final timepoint ($t=240$ hours) using binning values $\bm{x}^{\text{bin}}=\bm{y}^{\text{bin}}=\{0,40,80,\dots,1000\mu\text{m}\}$. We divide each histogram value by the binned area to estimate the spatial cell density. We vectorize each 2-dimensional density of size $25\times 25$ into a final vector of length 625.
2.2.3 Reduce dimension
We use principal components analysis (PCA) to reduce the dimension of each DV, resulting in dimensionality-reduced DVs (DRDVs). We use scree plots to display the explained variance of each principal component and determine a suitable dimension for the DRDVs [37]. We use the python package Scikit-learn (version 1.4.2) to perform PCA.
2.2.4 Data clustering
We use $k$-means clustering [38], an unsupervised machine learning method, to group each DRDV into one of $k$ clusters. The $k$-means algorithm clusters $d$-dimensional data by identifying $k$ mean vectors, $\bar{x}_{1},\bar{x}_{2},\dots,\bar{x}_{k}\in\mathbb{R}^{d}$. Each data sample is assigned into the cluster whose mean it is closest to, using Euclidean distance. More information on how the mean vectors are found is provided in [38]. We use the elbow method to determine the number of clusters to use for each dataset [37]. We use the python package Scikit-learn (version 1.4.2) to perform clustering.
2.2.5 Statistical Analysis
We perform a robustness assessment, visual inspection, and statistical tests to inspect and analyze the clustered data.
Robustness assessment: We test the robustness of the clustering assignments by computing out-of-sample (OOS) accuracy scores. We achieve this by labeling each parameter value according to which cluster most of its samples are placed into. A high (low) OOS accuracy score suggests that many simulations from the same parameter value will be placed into the same cluster (different clusters).
Computing OOS accuracy scores begins by splitting the 10 DRDVs from each parameter sample into 5 training and 5 testing DRDVs. We perform data clustering using the training DRDVs from all parameter values, and then label each parameter value according to the most common cluster assignments from its training DRDVs. To compute the OOS accuracy, we assign a “ground truth” label to each testing DRDV based on its parameter value’s label. We create a “predicted” label for each testing DRDV by running it through the trained clustering process. The OOS accuracy score is then given by the percentage of agreeing “ground truth” and “predicted” labels.
Visual inspection: Each cluster from Section 2.2.4 contains DRDVs that were created over many parameters values. We use visualization to better understand which parameter values generate each cluster. When varying a small number of parameters (such as 2 or 3), we inspect a parameter partition plot, which colors each parameter value according to its cluster labeling [34]. When varying a larger number of parameters, we create ridgeline plots for each cluster that depict the distribution of parameter samples within each cluster. Each parameter distribution is estimated using kernel density estimation [39]. We use the python package Joypy (version 0.2.4) to create ridgeline plots.
Statistical tests: For datasets generated by varying many parameters, we use the discrete two-sample Kolmogorov-Smirnov statistical test to evaluate whether the parameter distributions differ across clusters. For each varied parameter, we compare its samples between all (${k\choose 2}$ cluster pairs. Each comparison produces a $p$-value; low p-values suggest the parameter samples from the two clusters may be drawn from different distributions. We classify a parameter as impactful if the majority of these pairwise comparisons yield $p$-values below 0.05. We modified code from the ksdisc python package to perform statistical tests.
2.3 The Sobol’ GSA methodology
The Sobol’ Method is widely used to perform GSA [40, 41, 42]. We let $\bm{p}=(p_{1},p_{2},\dots,p_{Q})^{T}$ denote the input model parameters and $DV(\bm{p})$ represent the output DV summarizing an ABM simulation at the input parameter $\bm{p}$. The Sobol’ Method is a variance-based approach that estimates the contribution of each model parameter $p_{q}$ to the model’s total variance, $\text{Var}(DV(\bm{p}))$. Sobol’ introduced several sensitivity indices to quantify this contribution; we use the total-effect index in this study, which is given by [43, 44, 45]:
$S_{T_{q}}=\dfrac{\text{E}_{\bm{p}_{\sim q}}[\text{Var}_{p_{q}}(DV(\bm{p}_{\sim q}))]}{\text{Var}(DV(\bm{p}))},\ \ q=1,2,\dots,Q,$ (1)
where $\bm{p}_{\sim q}$ represents all parameters except for $p_{q}$, and $\text{E}_{\bm{p}_{\sim q}}(\cdot)$ and $\text{Var}_{\bm{p}_{\sim q}}(\cdot)$ represent the mean and variance over $\bm{p}$ while fixing $p_{q}$. This index quantifies the impact of changes to the input parameter $p_{q}$ on the model output $DV(\bm{p})$ while considering interactions between parameters.
The dimension of $S_{T_{q}}$ will match the dimension of $DV(\bm{p})$; suppose both are of length $m$. Our final sensitivity index is the averaged total effect index, which we calculate by averaging over all values in $S_{T_{q}}$:
$\bar{S}_{T_{q}}=\dfrac{1}{m}\sum_{i=1}^{m}(S_{T_{q}})_{i}.$ (2)
The values for $\bar{S}_{T_{q}}$ are typically estimated using quasi-Monte Carlo methods [46]. Saltelli [47, 48, 49] proposed an efficient sampling scheme to compute the total effect indices for all parameter values. This sampling requires $N(Q+2)$ parameter samples; it is common practice to choose $N$ to be a power of 2.
The Saltelli Dataset: As a comparison between the SSRCA and Sobol’ methods, we compute a third dataset generated using Saltelli parameter sampling [50]. Similar to Dataset 2, we generate this dataset by varying the same $Q=10$ parameters and choosing $N=2^{9}=512$ to ensure both datasets are of comparable size: Dataset 2 contains 5,445 samples while the Saltelli Dataset contains 6,144 samples. We implemented Saltelli sampling and the Sobol’ Index computations using the SALib package (version 1.5.1).
3 Results
3.1 Re-parameterizing the 2-dimensional Klowss Model
Computing the 3-dimensional ($\bm{x}=(x,y,z)$) Klowss Model across many parameter values is computationally prohibitive due to long simulation times. Instead, we simulate the 2-dimensional ($\bm{x}=(x,y)$) Klowss model, which corresponds to the equatorial cross section of the tumor spheroid at $z=0$. This reduction in dimension is justified by the radially symmetric nature of the model [21].
To account for this reduction, we modified two baseline parameter values: the number of starting agents, $N(0)$, and the ratio between nutrient consumption ($\kappa$) and diffusion ($D$) given by $\alpha=\kappa/D$. All other baseline values are set to those provided in [21]. To determine suitable $N(0)$ and $\alpha$ values, we simulated the 2-dimensional Klowss Model over a sample of both parameters and found that setting $N(0)=1100$ and $\alpha=0.01$ leads to close visual agreement with the $z=0$ cross section of the 3-dimensional Klowss Model (Supplementary Figures 7 and 8).
3.2 Summarizing the baseline model simulation
Our baseline simulation refers to the model simulation computed using the baseline parameter values from Table 1. This simulation begins with all red, yellow, and green cells at time 0 (Figure 3). An outer ring of green cells forms by day 2, and a necrotic core has developed by day 4. The ring and core are clearly established by day 10, and the spheroid has grown in size over the 10-day simulation.
Figure 3: Descriptor vectors (DVs) of the baseline ABM simulation (a-d) Snapshots at times t= 0, 2, 6, and 10 days from the baseline simulation. e) The Cell counts DV for the baseline simulation. f) The Final cell density DV for the baseline simulation. The solid (dashed) circle has a radius of 137 $\mu m$ (322 $\mu m$).
To summarize the baseline simulation, we compute its Cell counts and Final cell density DVs. According to the Cell counts DV, there is a steady increase in the number of red cells over the 10 day simulation, beginning with 850 and finishing with 3,000 cells, while the number of green and yellow cells remains relatively constant around 250 green cells and 10 yellow cells, respectively. The dead cell population initially grows slowly before increasing after $t=100$ hours; the simulation finishes with 1,000 dead cells. The Final cell density DV appears radially symmetric with an inner ring of 137 $\mu m$ and an outer ring of about 322 $\mu m$.
3.3 Analysis of Dataset 1
We explore the two-dimensional Klowss Model’s behavior when varying two parameters, $\eta_{1}$ and $c_{a}$, to generate Dataset 1. We summarized all 1,000 simulations in this dataset by computing the Cell counts and Final cell density DVs. We then reduced the dimensionality of all DVs using PCA and determined that 3-dimensional DRDVs are suitable representatives for the DVs (Supplementary Fig S9) for both DV types.
We trained the $k$-means clustering algorithm to the training datasets using both types of DRDVs separately and found that choosing $k=4$ or $5$ clusters can suitably cluster the data for both types of DRDVs. For simplicity, we present the results for $k=4$ clusters and note that the results are similar when setting $k=5$. When labeling each $(\eta_{1},c_{a})$ combination based on its most common clustering assignment, we find a clear partitioning of the $(\eta_{1},c_{a})$ parameter space into 4 distinct regions (Figure 4 and Supplementary Figures S11). The parameter partion plots that result from both types of DRDVs led to high OOS accuracy values, indicating that that these parameter designations are robust (Table 2).
Figure 4: Dataset 1 Analysis using the Cell counts DV. (a) The ($\eta_{1},c_{a}$) parameter partition plot for Dataset 1 when using the Cell counts DV. Snapshots at times $t=2,6,\text{ and }10$ days from the representative model simulations from (b-d) Cluster 1, (e-g) Cluster 2, (h-j) Cluster 3, and (k-m) Cluster 4.
Cell counts DRDVs Final cell density DRDVs
Dataset 1 99.2% 98.0%
Dataset 2 99.6% 98.1%
Table 2: Out of sample (OOS) accuracy scores for both datasets when using either the Cell counts or Final cell density DRDVs.
Investigation of the parameter partition plot for the Cell counts DRDVs suggests that $c_{a}$ is the most sensitive parameter, as each cluster contains only some of the considered $c_{a}$ parameter values but all considered $\eta_{1}$ values (Figure 4). Cluster 1 results for $c_{a}\leq 0.232$, Cluster 2 results for $0.232\leq c_{a}\leq 0.348$, Cluster 3 results for $0.306\leq c_{a}\leq 0.606$, and Cluster 4 results for $c_{a}\geq 0.459$. We visualized representative model simulations from each cluster; the simulation from Cluster 1 results in the earliest-forming and largest final necrotic core, whereas the simulation from Cluster 4 has the smallest necrotic core, which emerges around 8 days. Visual inspection of the Cell counts DVs of these representative simulations reveal that as we increase the cluster assignment from 1 to 4, the number of red cells decreases and the number of dead cells decreases (Supplementary Figure S13). Similarly, the Final cell density DVs show that the ring of living cells decreases in radius as the clustering number increases.
We did not perform statistical analysis between the clusters from this dataset due to the ease in visually interpreting each parameter’s impact in Figure 4 and Supplementary Figure S11.
3.4 Analysis of Dataset 2
We further explore the two-dimensional Klowss Model’s behavior by sampling all 10 “Varying” parameters from Table 1 to generate Dataset 2. We summarized all 54,450 simulations in this dataset by computing the Cell count and Final cell density DVs and found that 3-dimensional DRDVs are suitable representatives for the DVs (Supplementary Figure 9), and that $k=4$ clusters reliably cluster the data (Supplementary Figure 10). We trained the $k$-means clustering algorithm to the training datasets for both types of DRDVs, and the parameter labels that resulted from both types of DRDVs led to high OOS accuracy values (Table 2). The representative simulations from the four clusters demonstrate that Cluster 1 (4) simulations have the largest (smallest) necrotic cores. Representative DVs for these simulations quantify these differences (Supplementary Figure 14).
Ridgeline plots visualize the distribution of parameter samples within each cluster (Figure 5). The ridgeline plots when using Cell counts or Final cell density DRDVs are similar, so we present the Cell counts DRDV results here (Final cell density DRDV results shown in Supplementary Figure S15). Visual inspection of the distributions reveals that:
  • Cluster 1 emerges from low $c_{a}$ values;
  • Cluster 2 emerges from intermediate $c_{a}$ values, large $c_{d}$ values, low $\eta_{1}$ values, and low $\eta_{3}$ values;
  • Cluster 3 emerges from intermediate $c_{a}$ values, low to moderate $c_{d}$ values, and moderate to large $\eta_{1}$ values; and
  • Cluster 4 emerges from large $c_{a}$ values, and large $\eta_{1}$ values.
These observations suggest that the most sensitive models parameters are $c_{a}$, $c_{d}$, $\eta_{1}$, and $\eta_{3}$.
Figure 5: Dataset 2 Analysis using the Cell counts DV. (a-d) Ridgeline plots for Clusters 1-4. Each parameter $p$ is log-normalized as $\log_{2}(p/p_{base})$, where $p_{base}$ denotes the baseline model value from Table 1. Snapshots at times $t=2,6,\text{ and }10$ days from the representative model simulations from (e-g) Cluster 1, (h-j) Cluster 2, (k-m) Cluster 3, and (n-p) Cluster 4.
We identified sensitive model parameters by performing discrete two-sample Kolmogorov-Smirnov tests between all cluster pairs for all parameters when using the Cell counts DRDVs (Table 3). These tests confirm that $c_{a},c_{d},\eta_{1},\text{ and }\eta_{3}$ parameters are sensitive, and that all others parameters are insensitive. Performing the same sensitivity analysis using Final cell density DRDVs identifies these same four parameters are sensitive, as well as the $m_{min}$ parameter (Supplementary Table S5).
Cluster pairs
1 & 2 1 & 3 1 & 4 2 & 3 2 & 4 3 & 4 Classification
$c_{a}$ $<10^{-3}$ $<10^{-3}$ $<10^{-3}$ $<10^{-3}$ $<10^{-3}$ $<10^{-3}$ Sensitive
$c_{d}$ $<10^{-3}$ $<10^{-3}$ 0.059 $<10^{-3}$ $<10^{-3}$ $<10^{-3}$ Sensitive
$c_{m}$ 0.93 0.9 0.998 0.241 0.945 0.874 Insensitive
$d_{max}$ 0.34 0.916 1.0 0.963 0.303 0.893 Insensitive
$d_{min}$ 0.786 0.903 1.0 0.919 0.804 0.891 Insensitive
$m_{max}$ 0.659 0.892 0.999 0.956 0.653 0.898 Insensitive
$m_{min}$ 0.623 0.877 1.0 0.944 0.607 0.894 Insensitive
$\eta_{1}$ $<10^{-3}$ $<10^{-3}$ $<10^{-3}$ $<10^{-3}$ $<10^{-3}$ $<10^{-3}$ Sensitive
$\eta_{2}$ 0.929 0.87 1.0 0.843 0.93 0.891 Insensitive
$\eta_{3}$ $0.004$ $0.008$ 0.826 $<10^{-3}$ $<10^{-3}$ $<10^{-3}$ Sensitive
Table 3: Determining sensitive Dataset 2 parameters from the Cell counts DRDV. We designate $p$-values below 0.05 as being significant. If a parameter is significantly different for the majority of cluster pairs, then we classify the parameter as sensitive.
3.5 The Sobol’ and SSRCA Methods identify similar sensitive parameters
We compared the sensitivity analysis results that result from using the SSRCA and Sobol’ Methods. For the Sobol’ Method, we computed the averaged total effect effect indices, $\bar{S}_{T_{q}}$ from Equation (2), when varying all 10 parameters to generate the Saltelli Dataset. Both methods yield similar sensitivity analysis results. The $\bar{S}_{T_{q}}$ values are highest (indicative of high sensitivity) for the $c_{a}$ and $c_{d}$ parameters, followed by moderate $\bar{S}_{T_{q}}$ values for the $\eta_{1}$ and $\eta_{3}$ parameters; all other parameters have small $\bar{S}_{T_{q}}$ values (indicative of low sensitivity). The SSRCA method classifies these same four parameters as sensitive and all others as insensitive.
Figure 6: Comparing Sobol’ and SSRCA sensitivity analysis. We performed sensitivity analysis for the 10 model parameter using the Sobol’ indices methodology and the SSRCA methodology.
4 Discussion and conclusions
We introduce the Simulate, Summarize, Reduce dimensionality, Cluster, and statistically Analyze (SSRCA) Method as a new approach to perform sensitivity analysis (SA) for ABMs. The pipeline consists of summarizing ABM simulations with summary descriptor vectors (DVs) and then applying machine learning tools (dimensionality reduction and clustering) to these DVs to identify common model output patterns. The method concludes with statistical analysis between the clusters to interpret and identify which model parameter values most strongly influence these patterns. An important advantage of SSRCA over other common sensitivity analysis (SA) methods is its ability to simultaneously 1) determine sensitive parameters, 2) identify common model patterns, and 3) perform feature mapping for complex ABMs, even for a large number of parameters.
[…]
Morris Method Sobol’ Method SSRCA Method
Parameter sensitivity yes yes yes
Identify common yes no yes
Exhaustive feature mapping no no yes
Interaction effects no yes yes
Computationally efficient yes no further investigation required
Table 4: Comparing various capabilities of the Morris, Sobol’, and SSRCA Methods in performing sensitivity analysis.
Another strength of SSRCA is its flexibility with the choice of methodologies used within the pipeline. We found similar results in this work whether we used the Cell counts DV or Final Cell Density DV. This robustness to DV choice is crucial, as ABMs are broadly used for biological problems, and the appropriate DV choices depends on the application. For example, many studies rely on order parameters to characterize ABM behavior; alternatively, topological data analysis (TDA) is a recent field of research that has found success in summarizing many types of ABMs with applications in complex systems [36, 53, 54]. SSRCA can be used with any of these DVs, as it only requires they are uniform length. Additionally, we used principal components analysis and $k$-means clustering for dimensionality reduction and clustering, respectively, in this work, but any suitable algorithm can be used for these processes.
We applied the SSRCA methodology to a 2-dimsensional ABM of tumor spheroid growth and identified $c_{a}$, $c_{d}$, $\eta_{1}$, $\eta_{3}$ as the most sensitive parameters. This result suggests that the most impactful mechanisms for tumor spheroid phenotype are cell cycle entry and cell death, which may inform the future use of tumor spheroid models to interpret experimental data. For example, 3-dimsensional Klowss Model was calibrated to experimental data from human primary melanoma cells [21]; modeling similar data from other cell lines will likely require different parameter values. Our work proposes that thoroughly exploring all unknown parameters is not necessary; instead, one may more efficiently vary the 4 identified sensitive parameters to match the model to new data. Our proposed methodology may be of use for other modeling frameworks. For example, the authors of [55] were could not perform parameter estimation due to their differential equation model containing 19 unknown parameters. The SSRCA methodology could be applied to this model to learn an informative subset of its parameters for parameter estimation.
There are several avenues for future work with the SSRCA algorithm. While we can use the SSRCA Methodology to identify sensitive parameters, it is currently not able to rank parameters based on their sensitivity. We plan to further analyze the model and/or develop metrics that can be used in combination with SSRCA to enable such parameter ranking. We also demonstrated the utility of the SSRCA Method when applied to informative DVs, however, it has been reported previously that some DVs are more informative than others depending on the context [36, 34, 56, 57]. It would be interesting to explore if the SSRCA algorithm can be used to compare the information content between various DVS, or even to propose an optimal DV choice. We applied the SSRCA Methodology to a simple ABM of tumor spheroid growth in this work, but we plan to use it to interpret more complex models in the future, including models for other biological contexts, including cell biology, epidemiology, and ecology.
Data and code availability: All code and simulated data for this work is publicly available at
https://github.com/e-rohr/FUCCI_ABM2D.
Acknowledgements: The authors acknowledge use of the ELSA high performance computing cluster at The College of New Jersey for conducting the research reported in this paper. This cluster is funded in part by the National Science Foundation under grant numbers OAC-1826915 and OAC-2320244. The authors thank Matthew Simpson for useful discussion on this study.
Appendix A Model overview
We simulate an agent-based model of tumor spheroid growth developed in [21]. This model is hybrid, meaning we track both discrete agent locations and a continuous nutrient concentration (e.g., oxygen) within the spheroid. Tumor cells are represented as discrete agents centered at $\textbf{x}_{n}(t)=(x_{n}(t),y_{n}(t))$ for $n=1,\ldots,N(t)$, where $N(t)$ is the total number of agents at time $t$. The spatio-temporal distribution of the nutrient concentration is non-dimensionalized and represented by $c(\textbf{x},t)\in[0,1]$. In addition to agent locations, we also model how cells progress through the cell cycle. The agents are colored according to FUCCI: agents in the G1 phase are colored red, agents in the early S (eS) phase are colored yellow, and agents in the G2/S/M phase are colored green. Following [21], we assume the non-dimensional, local nutrient concentration, $c(\textbf{x},t)$, regulates cell motility, proliferation, and death [8, 58, 10]. The model is simulated in a cubic domain, $\Omega$, with side length $1000~{}\mu\mathrm{m}$, which we chose to mitigate computational costs while ensuring that agents do not reach the domain boundary during simulation.
We now describe how tumor cells in the model spheroid undergo mitosis, cell motility, and cell death, and then detail how the nutrient spreads and is consumed. All model parameters are described in Table 1. The tumor cell processes are modelled as discrete events using the Gillespie Algorithm [59]. We initialize the ABM in a similar manner as [21], where initial cell colors are randomized to produce proportions of red, yellow, and green cells that match experimental observations.
A.1 Agent Dynamics
We now discuss how agents progress through the cell-cycle. Each agent progresses at a rate determined by its cell cycle status. The red-to-yellow transition rate, $R_{r}(c)$, depends on the local nutrient concentration $c(\textbf{x},t),$ whereas the yellow-to-green transition rate, $R_{y}$, and the green-to-red transition rate, $R_{g}$, are independent of $c(\textbf{x},t)$. The the agent cell cycle progression rates are given by
$R_{r}(c)$ $=R_{r}\frac{c^{\eta_{1}}}{c_{a}^{\eta_{1}}+c^{\eta_{1}}},$ (3)
$R_{y}(c)$ $=R_{y},$ (4)
$R_{g}(c)$ $=R_{g},$ (5)
(See Table 1 for a description of all parameters). The cell-cycle transition from green to red includes mitosis, where a green agent produces two red agents. During this process, a polar angle $\theta$ is sampled to obtain a random direction where the daughter agents are placed. The first daughter agent is placed a distance $\mu/2$ in the randomly chosen direction, and the second daughter agent is placed $\mu/2$ units in the opposite direction, so that the two daughter cells are a distance $\mu$ apart The step length $\mu$ is chosen to be a typical cell diameter.
We now describe the model rules on agent migration and death. The agent migration and death rates, $m(c)$ and $d(c)$, respectively, are determined by the local nutrient concentration. Motility increases with $c$ since oxygen supports energy production; whereas, the death rate decreases with $c$ as cells are less likely to asphyxiate in high oxygen levels. These rates are given by
$m(c)$ $=(m_{max}-m_{min})\frac{c^{\eta_{2}}}{c_{m}^{\eta_{2}}+c^{\eta_{2}}}+m_{min},$ (6)
$d(c)$ $=(d_{max}-d_{min})\left(1-\frac{c^{\eta_{3}}}{c_{d}^{\eta_{3}}+c^{\eta_{3}}}\right)+d_{min},$ (7)
where all parameters are described in Table 1. When migrating, an agent is displaced $\mu$ units in the randomly chosen direction. When an agent dies, it is removed from the simulation by setting its status to dead, and its final location is recorded.
A.2 Nutrient dynamics
The spatio-temporal distribution of the dimensionalized nutrient concentration, $C(\textbf{x},t)$, is assumed to be governed by the reaction-diffusion equation
$\frac{\partial C}{\partial t}=D\nabla^{2}C-\kappa Cv,\text{ in }\Omega,$ (8)
with diffusivity $D>0~{}[\mu\mathrm{m}^{2}~{}\mathrm{h}^{-1}]$, consumption rate $\kappa>0~{}[\mu m^{2}~{}(\mathrm{h}~{}\text{cells})^{-1}]$, and where $v(\textbf{x},t)\geq 0~{}[\text{cells }\mu m^{-2}]$ is the density of agents at position x and time $t.$ On the boundary, $\partial\Omega$, a maximum far-field concentration $C(\textbf{x},t)=c_{b}$ is imposed. We non-dimensionalize this independent variable by setting $c(\textbf{x},t)=C(\textbf{x},t)/c_{b}$, so that $c=1$ on $\partial\Omega$, and $c(\textbf{x},t)\in[0,1]$. We approximate the reaction-diffusion Equation (8) using the quasi-steady state equation
$0=\nabla^{2}c-\alpha cv,\text{ in }\Omega,$ (9)
where $\alpha=\kappa/D>0\ [\mu\mathrm{m}~{}\mathrm{cell}^{-1}]$. During simulations, the spatio-temporal distribution of the chemical is modelled by solving Equation (9) $t^{*}$ units using a finite volume method on a uniform structured mesh with node spacing $h$. Following [21], we use $t^{*}=1$. The cell density $v(\textbf{x},t)$ is approximated by setting $v(\textbf{x}_{i,j},t)=N_{i,j}/h^{2}$, where $N_{i,j}$ is the number of agents within the control volume surrounding the node located at $(x_{i},y_{j})$ and $h^{2}$ is the volume of the control volume.
Figure 7: Determining the baseline $N(0)$ value. To determine the baseline value of $N(0)$, we visually compared simulations of the 2-dimensional Klowss model (top 5 rows) for various values of $N(0)$ to the baseline simulation of the 3-dimensional Klowss Model from [21] (bottom row). The latter depicts the $z=0$ cross section of the 3-dimensional simulation. The red text label indicates the parameter value we determined most closely matches the bottom row. The bottom row is reproduced and modified from [21], Copyright Klowss et al., and licensed under CC-BY 4.0 https://creativecommons.org/licenses/by/4.0/.
Figure 8: Determining the baseline $\alpha$ value. To determine the baseline value of $\alpha$, we visually compared simulations of the 2-dimensional Klowss model (top 5 rows) for various values of $\alpha$ to the baseline simulation of the 3-dimensional Klowss Model from [21] (bottom row). The latter depicts the $z=0$ cross section of the 3-dimensional simulation. The red text label indicates the parameter value we determined most closely matches the bottom row. The bottom row is reproduced and modified from [21], Copyright Klowss et al., and licensed under CC-BY 4.0 https://creativecommons.org/licenses/by/4.0/.
Figure 9: Scree plots. The scree plots that result when performing dimensionality reduction using (a) Dataset 1 using the Cell counts DRDV, (b) Dataset 1 using the Final cell density DRDV, (c) Dataset 2 using the Cell counts DRDV, (d) Dataset 2 using the Final cell density DRDV
Figure 10: Elbow plots. The elbow plots that result when clustering data from (a) Dataset 1 using the Cell counts DRDV, (b) Dataset 1 using the Final cell density DRDV, (c) Dataset 2 using the Cell counts DRDV, (d) Dataset 2 using the Final cell density DRDV
Figure 11: Dataset 1 Analysis using the Final cell density DV. (a) The ($\eta_{1},c_{a}$) Parameter partition plot for Dataset 1 when using the Cell counts DV. Snapshots at times $t=2,6,\text{ and }10$ days from the representative model simulations from (b-d) Cluster 1, (e-g) Cluster 2, (h-j) Cluster 3, and (k-m) Cluster 4.
Figure 12: Confusion matrices. The confusion matrices that result when performing clustering robustness assessment for (a) Dataset 1 using the Cell counts DRDV, (b) Dataset 1 using the Final cell density DRDV, (c) Dataset 2 using the Cell counts DRDV, (d) Dataset 2 using the Final cell density DRDV
Figure 13: Represenative DVs for Dataset1. Representative Cell counts DVs for Dataset 1 from (a-d) clusters 1-4. Representative Final cell density DVs for Dataset 1 from (e-h) clusters 1-4.
Figure 14: Represenative DVs for Dataset2. Representative Cell counts DVs for Dataset 2 from (a-d) clusters 1-4. Representative Final cell density DVs for Dataset 2 from (e-h) clusters 1-4.
Figure 15: Dataset 2 Analysis using the Final cell density DV. (a-d) Ridgeline plots for Clusters 1-4. Each parameter $p$ is log-normalized as $\log_{2}(p/p_{base})$, where $p_{base}$ denotes the baseline model value from Table 1. Snapshots at times $t=2,6,\text{ and }10$ days from the representative model simulations from (e-g) Cluster 1, (h-j) Cluster 2, (k-m) Cluster 3, and (n-p) Cluster 4.
Cluster pairs
1 & 2 1 & 3 1 & 4 2 & 3 2 & 4 3 & 4 Classification
$c_{a}$ $<10^{-3}$ $<10^{-3}$ $<10^{-3}$ $<10^{-3}$ $<10^{-3}$ $<10^{-3}$ Sensitive
$c_{d}$ $<10^{-3}$ $0.008$ $<10^{-3}$ $<10^{-3}$ $<10^{-3}$ $<10^{-3}$ Sensitive
$c_{m}$ 0.307 0.626 0.995 $<10^{-3}$ 0.098 0.712 Insensitive
$d_{max}$ $0.048$ 0.915 0.999 1.0 $0.027$ 0.911 Insensitive
$d_{min}$ 0.767 0.514 0.883 0.525 0.961 0.875 Insensitive
$m_{max}$ $0.023$ 0.884 0.999 0.998 $0.016$ 0.9 Insensitive
$m_{min}$ $<10^{-3}$ 0.964 1.0 $<10^{-3}$ $<10^{-3}$ 0.911 Sensitive
$\eta_{1}$ $0.048$ 0.061 $<10^{-3}$ 0.215 $0.01$ 0.246 Sensitive
$\eta_{2}$ 0.934 0.784 0.998 0.361 0.975 0.878 Insensitive
$\eta_{3}$ $<10^{-3}$ $<10^{-3}$ 1.0 $<10^{-3}$ $<10^{-3}$ $<10^{-3}$ Sensitive
Table 5: Determining sensitive Dataset 2 parameters from the Final cell density DRDV. We designate $p$-values below 0.05 as being significant. If a parameter receives $p$-values below 0.05 for the majority of cluster pairs, then we classify the parameter as sensitive.